solve:
a^2-6a-27=0
help?

Found 2 solutions by ewatrrr, jessica43:
Answer by ewatrrr(24785)


Hi
solving
a^2-6a-27=0
factoring
(a-9)(a+3)= 0 Note:SUM of the inner product(-9a) and the outer product(3a) = -6a
(a-9)=0
a = 9
(a+3)= a = -3

Answer by jessica43(140)

x^2-6x-27=0
To solve this problem you need to factor the trinomial, which is the opposite of the FOIL method. The factors will have this form:
(x A)(x B)
where A and B will be the factors of the third term (-27). The factors of -27 are: -1, 1, -3, 3,-9, 9, -27, and 27. So these are our possibilities of A and B: -1&27, 1&-27, -3&9, 3&-9.
Now, we need to see which of these combinations, when added, will equal -6 (since the second term of our trinomial is -6x).
The only combination that could equal -6 when added is -9 and 3.
So:
x^2 - 6x - 27 = 0
(x-9)(x+3)= 0
So for this to equal zero, x = 9 or x = -3.
Your answers are x = 9 and x = -3.